# DCPI-Depth: Explicitly Infusing Dense Correspondence Prior to Unsupervised Monocular Depth Estimation

### v1

@@ III-D Bidirectional Stream Co-Adjustment Strategy
Figure 4: Qualitative comparison among SC-DepthV1, SC-DepthV3, and our DCPI-Depth on the KITTI [45] dataset.
IV Experiments
IV-A Datasets
The KITTI[45] dataset contains 56 different scenes, evenly divided for model training and evaluation. Well-rectified stereo image pairs with depth information acquired from the LiDAR point clouds are provided. In our experiments, we adopt the Eigen split [5], which comprises 39,180 monocular triplets for training, 4,424 images for validation, and 697 images for testing.
The DDAD[31] dataset consists of 200 urban driving scene sequences, with LiDAR point clouds provided to generate sparse ground-truth depth for model performance evaluation. Following [13], we split the dataset into a training set of 12,650 images and a test set of 3,950 images. As optical flow ground truth is unavailable in this dataset, we use FlowNet pre-trained on the KITTI dataset to provide dense correspondence priors.
The Make3D[52] dataset contains 134 images of outdoor scenes along with their corresponding depth maps. Since neither stereo image pairs nor monocular sequences are provided, we cannot use this dataset to train monocular depth estimation networks in a unsupervised manner. Therefore, we only use this dataset to quantify the generalizability of models per-trained on the KITTI dataset.
IV-B Experimental Setup
Our experiments are conducted on an NVIDIA RTX 4090 GPU with a batch size of 12. Following [10], we adopt a training approach wherein a snippet of three consecutive video frames is utilized as a training sample. To augment the dataset, random color jitter and horizontal flips are applied to the images during model training. We use AdamW [53] optimizer to minimize the loss functions, and train the network for 30 epochs with an initial learning rate of $10^{-4}$ and a weight decay of $10^{-2}$. Following [10, 14], the network’s encoder is initialized using pre-trained weights from the ImageNet database [54]. RAFT [17], a highly efficient optical flow estimation network pre-trained on the KITTI Flow 2015 dataset, is employed as the FlowNet in our framework to provide dense correspondence prior.
IV-C Evaluation Metrics
@@
TABLE IV: Ablation studies on the KITTI [45] dataset. The best results are shown in bold symbol. The symbols $\uparrow$ and $\downarrow$ indicate that higher and lower values correspond to better performance, respectively.
DepthNet Backbone CPG Stream BSCA Strategy Full Image Static Regions Dynamic Regions
CGDC loss DPC loss Abs Rel $\downarrow$ Sq Rel $\downarrow$ RMSE $\downarrow$ $\delta_{1}$uparrow $\delta_{2}$uparrow Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\delta_{1}$uparrow Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\delta_{1}$uparrow
ResNet18 [55] Baseline 0.118 0.756 4.709 0.864 0.960 0.108 0.636 0.881 0.205 2.283 0.703
0.111 0.698 4.584 0.874 0.963 0.103 0.601 0.888 0.193 2.125 0.718
0.115 0.716 4.611 0.869 0.963 0.105 0.610 0.886 0.201 2.234 0.710
0.113 0.707 4.559 0.873 0.964 0.103 0.591 0.891 0.211 2.510 0.700
0.109 0.679 4.496 0.878 0.965 0.101 0.584 0.892 0.186 1.948 0.732
DIFFNet [8] Baseline 0.108 0.696 4.591 0.878 0.964 0.099 0.603 0.892 0.197 2.188 0.717
0.103 0.653 4.483 0.886 0.966 0.094 0.555 0.900 0.191 2.098 0.719
IV-D Comparison with SoTA Approaches
The quantitative experimental results presented in Table I demonstrate that DCPI-Depth achieves SoTA performance when processing images with resolutions of 192$\times$640 pixels and 320$\times$1024 pixels. Specifically, our DCPI-Depth outperforms all other SoTA un/self-supervised methods, achieving a 4.0% reduction in Abs Rel and a 0.6% improvement in $\delta_{1}$ at a resolution is 192$\times$640 pixels. Given the known limitations of the ground truth quality in the KITTI dataset, we further evaluate our network performance using the KITTI improved ground truth [50]. It can be observed in Table II that DCPI-Depth consistently outperforms the previous SoTA method [35] across all metrics.
The qualitative experimental results on the KITTI dataset are shown in Fig. 3. Our DCPI-Depth exhibits superior performance compared to previous SoTA methods. This is particularly evident in texture-less regions, such as (a) and (b). Additionally, our approach ensures smoother and more continuous depth changes and generates clearer boundaries, as exemplified in (c) to (f). These improvements are attributed to the dense correspondence priors infused via our proposed CGDC and DPC losses within the CPG stream, which provide refined and direct geometric cues for depth estimation. Furthermore, DCPI-Depth maintains accurate depth estimation on dynamic objects, such as (g) and (h), unaffected by independent motion. This advantage, observed on dynamic objects, stems from our BSCA strategy, which not only prevents the misleading effects of independent motion captured by the optical flow within the CPG stream but also reinforces accurate estimations early in the training phase, preserving them through to the final results.
@@ IV-D Comparison with SoTA Approaches
Figure 7: Qualitative zero-shot monocular depth estimation results on the Make3D [52] dataset.
To further validate the effectiveness of DCPI-Depth in leveraging dense correspondence priors, we deploy our framework to SC-DepthV3 [13], which leverages pseudo depth to achieve robust and highly reliable depth estimation. It also provides a comprehensive platform for evaluating depth performance across dynamic objects, static areas, and full images. As shown in Table III, DCPI-Depth achieves significant improvements, outperforming the baseline by a considerable margin across all metrics for full image, static regions, and dynamic regions. Furthermore, we provide comparisons of learning curves among SC-DepthV1, SC-DepthV3, and DCPI-Depth. As illustrated in Fig. 5(a), DCPI-Depth significantly improves depth estimation performance compared to SC-DepthV3. These results demonstrate that the dense correspondence prior we incorporate additionally provides meaningful geometric cues on top of pseudo depth. Moreover, it can be observed in Fig. 5(b) that the convergence of the photometric loss is comparable among the three models. This observation suggests that merely minimizing photometric loss, which provides an indirect supervisory signal, presents challenges in further improving depth estimation performance. In contrast, our approach provides a more direct and effective constraint for depth estimation. The qualitative results shown in Fig. 4 also indicate that DCPI-Depth excels, particularly in preserving clear object boundaries and maintaining robustness in ambiguous regions. This validates the effectiveness of incorporating dense correspondence priors into depth estimation frameworks.
@@ IV-E Ablation Studies
Table IV and V presents comprehensive ablation studies conducted with SC-DepthV3 to validate the effectiveness of our contributed components. The first ablation study validates the internal design of the CPG stream by comparing the overall performance with and without the incorporation of the CGDC loss and DPC loss, respectively. Key findings from this study include: (1) employing either the CGDC loss or the DPC loss independently leads to improved performance; (2) the combined use of both losses leads to more substantial improvements compared to using each loss individually.
@@
Finally, as shown in Table V, we evaluate the performance of our approach with respect to different qualities of dense correspondence priors. These results suggest the robustness of our method, demonstrating effective performance even when the dense correspondences are of low quality (when FlowNet is trained with an insufficient number of iterations). We attribute this phenomenon to the fact that even low-quality dense correspondence priors can provide valuable guidance for training DepthNet, helping perceive relative distances among pixels and reflecting depth variations to some extent. Meanwhile, this also benefits from our BSCA strategy, which effectively allows the two types of flows to complement each other.
TABLE V: Ablation study to demonstrate the effectiveness of our approach with respect to different dense correspondence prior qualities on the KITTI [45] dataset. “F1-EPE” (pixels) and “F1-All” (pixels) are two metrics used to quantify the performance of FlowNet, where lower values indicate better performance.
       Pre-trained FlowNet Optical Flow Quality Depth Estimation Metrics
F1-EPE F1-All Abs Rel Sq Rel RMSE $\delta_{1}$ $\delta_{2}$
      w/ 0.63 1.50 0.109 0.679 4.496 0.878 0.965
      w/ 1.94 9.83 0.113 0.695 4.640 0.871 0.963
      w/o - - 0.118 0.756 4.709 0.864 0.960
TABLE VI: Quantitative experimental results for SoTA monocular depth estimation models, with and without our proposed DCPI-Depth framework employed, using the KITTI Eigen split [5]. All models are trained on 192$\times$640 images.
@@
Lite-Mono-8M[14] 0.101 0.729 4.454 0.178 0.897 0.965 0.983
+Ours 0.095 0.662 4.274 0.170 0.902 0.967 0.985
TABLE VII: Quantitative results on the DDAD [31] dataset.
Method Abs Rel Sq Rel RMSE RMSE log $\delta_{1}$ $\delta_{2}$ $\delta_{3}$
SC-DepthV1 0.169 3.877 16.290 0.280 0.773 0.905 0.951
SC-DepthV3 0.142 3.031 15.868 0.248 0.813 0.922 0.963
DCPI-Depth (Ours) 0.143 2.963 15.380 0.244 0.812 0.925 0.966
TABLE VIII: Quantitative results on the Make3D [52] dataset. All models are trained on the KITTI [45] dataset images (resolution: 640$\times$192 pixels).
Method Abs Rel Sq Rel RMSE RMSE log
Monodepth2 [10] 0.321 3.378 7.252 0.163
@@ IV-F Generalizability Evaluation
We incorporate the CPG stream and the BSCA strategy into existing open-source methods to further demonstrate the adaptability of our contributions to other SoTA methods. As shown in Table VI, our proposed stream and strategy significantly improve the performance of the original networks across all metrics, which consistently validates the scalability of our methods, as detailed in Sect. IV-E.
In addition, we employ FlowNet, pre-trained on the KITTI Flow 2015 [56] dataset, within our DCPI-Depth framework, and train our model on the DDAD [31] dataset to evaluate the generalizability of the dense correspondence prior provided by the pre-trained FlowNet. The results presented in Table VII indicate that our method improves the depth estimation performance across five metrics. However, the improvements are not as significant as those observed when trained on the KITTI dataset. This limitation may be due to the domain gap between these two datasets.
Finally, we conduct a zero-shot experiment on the Make3D [52] dataset using the pre-trained weights from KITTI. As shown in Table VIII and Fig. 7, our model outperforms all other methods, demonstrating its ability to generalize effectively to new, unseen scenes.
@@ V Conclusion
This article presented DCPI-Depth, a novel unsupervised monocular depth estimation framework with two bidirectional and collaborative streams: a conventional PCG stream and a newly developed CPG stream. The latter was designed specifically to infuse dense correspondence priors into monocular depth estimation. It consists of a CGDC loss to provide contextual-based depth with geometric guidance obtained from ego-motion and dense correspondence priors, and a DPC loss to constrain the local depth variation using the explicit relationship between the differential properties of depth and optical flow. Moreover, a BSCA strategy was developed to enhance the interaction between the two flow types, encouraging the rigid flow towards more accurate correspondence and making the optical flow more adaptable across various scenarios under the static scene hypotheses. Compared to previous works, our DCPI-Depth framework has demonstrated impressive performance and superior generalizability on the KITTI, DDAD, and Make3D datasets. Future work will focus on exploring more advanced optical flow estimation techniques to enhance the reliability and generalizability of dense correspondence priors and to extend the techniques in DCPI-Depth into a joint learning framework where multiple tasks can explicitly couple and mutually enhance each other.

### v2

@@ IV-C Evaluation Metrics
MonoDiffusion[47] 2024 320 $\times$ 1024 M 0.095 0.670 4.219 0.171 0.909 0.968 0.984
IV-D Comparison with SoTA Approaches
The quantitative experimental results presented in Table I demonstrate that DCPI-Depth achieves SoTA performance across the KITTI [43], DDAD [29], nuScenes [44], and Waymo Open [45] datasets. Notably, the employed baseline model, Lite-mono [12], that performs unsatisfactorily on these datasets, achieves significant performance improvements, with error reductions ranging from 13.04% to 61.81% on Abs Rel. This substantial enhancement enables DCPI-Depth to surpass previously leading methods [11, 42] on each of the respective datasets, suggesting the effectiveness of our proposed framework.
The qualitative experimental results on the KITTI dataset are shown in Fig. 4. Our DCPI-Depth exhibits superior performance compared to previous SoTA methods. This is particularly evident in texture-less regions, such as (a) and (b). Additionally, our approach ensures smoother and more continuous depth changes and generates clearer boundaries, as exemplified in (c) to (f). These improvements are attributed to the dense correspondence priors infused via our proposed CGDC and DPC losses within the CPG stream, which provide refined and direct geometric cues for depth estimation. Furthermore, DCPI-Depth maintains accurate depth estimation on dynamic objects, such as (g) and (h), unaffected by independent motion. This advantage, observed on dynamic objects, stems from our BSCA strategy, which not only prevents the misleading effects of independent motion captured by the optical flow within the CPG stream but also reinforces accurate estimations early in the training phase, preserving them through to the final results. As shown in Fig. 5, the baseline model performs poorly on the DDAD, nuScenes, and Waymo datasets, but demonstrates dramatic performance improvements after trained under our proposed framework, further demonstrating the robustness of DCPI-Depth across diverse scenarios.
To further investigate the impact of image resolutions and ground truth quality, We evaluate our network’s performance using (1) images at a resolution of 320$\times$1024 pixels and (2) the improved ground truth [53] from the KITTI dataset. It can be observed in Table II that DCPI-Depth consistently achieves SoTA performances compared with existing methods. Although certain metrics exhibit suboptimal results, potentially attributable to the inherent limitation of the network backbone, the proposed training framework, infused with dense correspondence priors, demonstrates a marked improvement in baseline performance and robustness.
@@ IV-D Comparison with SoTA Approaches
TABLE IV: Ablation studies on the KITTI [43] and the DDAD [29] datasets. The best results are shown in bold symbol. The symbols $\uparrow$ and $\downarrow$ indicate that higher and lower values correspond to better performance, respectively.
DepthNet Backbone CPG Stream BSCA Strategy KITTI (Full Image) KITTI (Dynamic Regions) DDAD (Full Image)
CGDC loss DPC loss Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\delta_{1}$uparrow $\delta_{2}$uparrow Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\delta_{1}$uparrow Abs Rel $\downarrow$ Sq Rel $\downarrow$ $\delta_{1}$uparrow $\delta_{2}$uparrow
ResNet18 [54] Baseline 0.118 0.756 0.864 0.960 0.205 2.283 0.703 0.149 3.062 0.798 0.920
0.111 0.698 0.874 0.963 0.193 2.125 0.718 0.143 3.156 0.801 0.915
0.115 0.716 0.869 0.963 0.201 2.234 0.710 0.148 3.092 0.799 0.921
0.113 0.707 0.873 0.964 0.211 2.510 0.700 0.145 3.066 0.805 0.919
0.109 0.679 0.878 0.965 0.186 1.948 0.732 0.143 2.963 0.812 0.925
DIFFNet [7] Baseline 0.108 0.696 0.878 0.964 0.197 2.188 0.717 0.145 3.104 0.804 0.922
0.103 0.653 0.886 0.966 0.191 2.098 0.719 0.137 2.812 0.816 0.926
Figure 7: Qualitative results on the KITTI [43] dataset to demonstrate the effectiveness of the CGDC loss and DPC loss.
@@
To further validate the effectiveness of DCPI-Depth in leveraging dense correspondence priors, we deploy our framework to SC-DepthV3 [11], which leverages pseudo depth to achieve robust and highly reliable depth estimation. It also provides a comprehensive platform for evaluating depth performance across dynamic objects, static areas, and full images. As shown in Table III, DCPI-Depth achieves significant improvements, outperforming the baseline by a considerable margin across all metrics for full image, static regions, and dynamic regions. Furthermore, we provide comparisons of learning curves among SC-DepthV1, SC-DepthV3, and DCPI-Depth. As illustrated in Fig. 6(a), DCPI-Depth significantly improves depth estimation performance compared to SC-DepthV3. These results demonstrate that the dense correspondence prior we incorporate additionally provides meaningful geometric cues on top of pseudo depth. Moreover, it can be observed in Fig. 6(b) that the convergence of the photometric loss is comparable among the three models. This observation suggests that merely minimizing photometric loss, which provides an indirect supervisory signal, presents challenges in further improving depth estimation performance. In contrast, our approach provides a more direct and effective constraint for depth estimation.
@@ IV-E Ablation Studies
Table IV presents comprehensive ablation studies conducted with SC-DepthV3 to validate the effectiveness of our contributed components. The first ablation study validates the internal design of the CPG stream by comparing the overall performance with and without the incorporation of the CGDC loss and DPC loss, respectively. Key findings from this study include: (1) Incorporating the CGDC loss results in the most significant performance improvements. (2) The DPC loss also plays an important role in improving the performance of our framework. Without the DPC loss, the Abs Rel metric, which measures the relative depth error similar to the CGDC loss, remains steady. However, other metrics decrease significantly, especially on the DDAD dataset. Fig. 7 provides quantitative experimental results to demonstrate the effectiveness of the CGDC and DPC losses. It is obvious that removing the CGDC loss can lead to significant depth distribution shifts (see the first and third lines), which aligns with our expectations. Furthermore, despite achieving satisfactory quantitative results, removing the DPC loss can result in erroneous depth estimations in continuous regions such as vehicles and cyclists, as illustrated in the second to fourth lines.
@@
TABLE V: Quantitative experimental results for additional monocular depth estimation models, with and without our proposed DCPI-Depth framework employed, using the KITTI Eigen split [5]. All models are trained on 192$\times$640 images.
@@
TABLE VI: Quantitative results on the Make3D [49] and DIML [50] datasets. All models are trained on the KITTI [43] dataset images (resolution: 640$\times$192 pixels).
Dataset Method Abs Rel Sq Rel RMSE RMSE log
Make3D Monodepth2 [9] 0.321 3.378 7.252 0.163
@@
DIML Monodepth2 [9] 0.185 0.298 1.140 0.249
HR-Depth [39] 0.183 0.296 1.128 0.248
R-MSFM6 [56] 0.181 0.301 1.132 0.243
Lite-Mono [12] 0.173 0.271 1.108 0.239
MonoDiffusion [47] 0.166 0.256 1.084 0.232
DCPI-Depth (Ours) 0.163 0.237 1.038 0.226
@@ IV-F Generalizability Evaluation
We incorporate the CPG stream and the BSCA strategy into existing open-source methods to further demonstrate the adaptability of our contributions to other SoTA methods. As shown in Table V, our proposed stream and strategy significantly improve the performance of the original networks across all metrics, which consistently validates the scalability of our methods, as detailed in Sect. IV-E.
Finally, we conduct zero-shot experiments on the Make3D [49] and DIML [50] datasets using the models pretrained on the KITTI dataset. As shown in Table VI, our model outperforms all other methods across these two datasets, demonstrating its ability to generalize to new, unseen scenes.
@@ V Conclusion
This article presented DCPI-Depth, a novel unsupervised monocular depth estimation framework with two bidirectional and collaborative streams: a conventional PCG stream and a newly developed CPG stream. The latter was designed specifically to infuse dense correspondence priors into monocular depth estimation. It consists of a CGDC loss to provide contextual-based depth with geometric guidance obtained from ego-motion and dense correspondence priors, and a DPC loss to constrain the local depth variation using the explicit relationship between the differential properties of depth and optical flow. Moreover, a BSCA strategy was developed to enhance the interaction between the two flow types, encouraging the rigid flow towards more accurate correspondence and making the optical flow more adaptable across various scenarios under the static scene hypotheses. Compared to previous works, our DCPI-Depth framework has demonstrated impressive performance and superior generalizability across six public datasets. Future work will focus on exploring more advanced optical flow estimation techniques to enhance the reliability and generalizability of dense correspondence priors and to extend the techniques in DCPI-Depth into a joint learning framework where multiple tasks can explicitly couple and mutually enhance each other.